OpenStudy (anonymous):

(8-3sqrt6)(7+2sqrt6), solve this one too plz,

OpenStudy (anonymous):

Apply foil to the problem with the following result: \[56-36-21\sqrt{6}+16\sqrt{6}=20-5\sqrt{6}\]

OpenStudy (anonymous):

oooh, i missed a step thats why it was wrong, thx for corecting me